# Operator Deep Smoothing for Implied Volatility

### v2

@@ 3.2 Operator Deep Smoothing
(a) Smoothed surface with input volatility quotes (blue). $\delta_{\text{abs}}=0.004$, $\delta_{\text{spr}}(\hat{v},\mathbf{v})=1.061$.
(b) Slices of implied volatility (left) and implied density (right), scaled by time-to-expiry. Left: Absence of crossings indicates absence of calendar arbitrage. Right: Positivity indicates absence of butterfly arbitrage. In fact $\mathcal{L}_{\text{cal}}(\theta,\mathbf{v})=6.3\times 10^{-6}$, $\mathcal{L}_{\text{but}}(\theta,\mathbf{v})=0$.
Figure 2: Operator deep smoothing of quotes $\mathbf{v}$ from 04.01.2021 at 10:50:00. Compare Figure 10.
Let $\mathbf{v}=\{v(x_{l})\}_{l=1}^{p}$ be the collection of observed data, for example implied volatilities as in Section 2. This notation silently adopts the neural operator philosophy, connecting the data point $v(x_{l})$ with coordinates $x_{l}$, hinting at a latent function $v\colon D\to\mathbb{R}$ giving rise to the observed values. The smoothing or interpolation task consists of constructing an appropriately regular candidate $\hat{v}$ for $v$ from the data $\mathbf{v}=v|_{\pi}$, with $\pi=\{x_{1},\dots,x_{p}\}$. The operator deep smoothing approach uses a neural operator $\tilde{F}^{\theta}$, trained using historical data, to generate $\hat{v}:=\tilde{F}^{\theta}(\mathbf{v})$. It fundamentally leverages the discretization-invariance to produce consistent results even if the ”sensors” $x_{l}$ of the latent function $v$ continuously change in availability and/or location in the domain $D$. This is the situation for volatility smoothing, where the sensors $x_{l}=(\tau_{l},k_{l})$ are the (time-to-expiry, log-moneyness) coordinates of the quoted options, which – as noted in Section 1 and illustrated in Figure 1(b) – move with the market, thus motivating the operator deep smoothing approach. We now describe the methodology in more detail. For a first glance at the results we refer to Figure 2.
@@
Translating the smoothing task to an operator learning problem, we find that the target operator is the (continuous) identity operator $F=F_{\text{id}}\colon\mathcal{A}\ni v\mapsto v\in\mathcal{U}$, with equal input and output spaces $\mathcal{A}=\mathcal{U}=C(\overline{D})$. The training dataset is the collection $\mathcal{D}=\{\mathbf{v}^{(i)}\}$ of historical data (labels and features coincide), and we train a suitable neural operator architecture $\tilde{F}^{\theta}$ such that
$|\tilde{F}^{\theta}(\mathbf{v})(x)-v(x)|\leq\varepsilon,\qquad\text{for all }x\in D,$ (7)
for some given error tolerance $\varepsilon$ for the interpolation task, while $v$ is the latent function of which we observe $\mathbf{v}=v|_{\pi}$. Instead of minimizing empirical estimates of the $\infty$-norm $\|\tilde{F}^{\theta}(\mathbf{v})-v\|_{\mathcal{U}}$, however, we suggest a fitting loss based on the root mean square relative error,
@@
for its smoothness and invariance to the scale of the data. [^4] Empirical estimates of $\infty$-norms and $L^{2}$-norms are equivalent loss functions on finite-dimensional spaces. Additional engineering techniques, such as sub-sampling of inputs during training, are explored in our practical investigation in Section 4.
Practical Constraints
Depending on the application, the smoothing task may be subject to constraints. For volatility smoothing, the smoothed surface $\hat{v}^{\theta}=\tilde{F}^{\theta}(\mathbf{v})$ must be free of static arbitrage. This is effectively enforced by augmenting the loss function with additional penalization terms, moving away from a pure operator learning problem. [^5] It is not the goal of volatility smoothing to learn the identity operator throughout the entire input space. This does not only promote the relevant properties in the neural operator output but can also help define it when faced with sparsity of data in the domain $D$ (in this context see also Li et al., (2021)). From Theorem 2.1, these penalization terms naturally motivate our choices of $\mathcal{L}_{\text{but}}$ and $\mathcal{L}_{\text{cal}}$. The strike arbitrage constraint 2.1(i) is handled via
@@
where we ignore the asymptotic condition since our experiments are focused on the bounded domain $D$ (Figure 1(a)). The inclusion of $\varepsilon$ promotes strictly positive implied densities (we will use $\varepsilon=10^{-3}$), while we choose the $1$-norm to induce sparsity in the constraint violation. The calendar arbitrage constraint 2.1(ii) can be tackled analogously with
@@
Various neural operator architectures exist, mostly arising from the kernel integral transform framework of Kovachki et al., (2023). Most prominently, these include Fourier neural operators (FNO), delivering state-of-the-art results on fixed grid data, as well as graph neural operators (GNO), able to handle arbitrary mesh geometries (both reviewed in Appendix A.1). While highly effective with documented universality, these neural operators are not directly applicable for interpolation tasks as their layers include a pointwise-applied linear transformation, which limits the output to the set of the input data locations. Dropping this local transformation results in an architecture proved to retain universality Kovachki et al., (2023) and that – at least for its implementation as a GNO – allows to interpolate functions. On the other hand, Lanthaler et al., (2023) proves universality for the architecture combining the local linear transformation with a simple averaging operation, suggesting the fundamental importance of the collaboration of local and non-local components for the expressivity of neural operators. This was noted in Kovachki et al., (2023), for whom retaining the local components can be “beneficial in practice”, and confirmed in our experiments, where a purely non-local architecture led to substantially reduced performance.
@@ 4.1 Model Training
We perform our numerical experiments using 20-minute intervals of CBOE S&P 500 Index Option data from 2012 to 2021. The dataset amounts to a collection of 49089 implied volatility surfaces and just above 60 million individual volatility datapoints (after domain truncation). We refer the reader to Appendix C.1 for full details on the preparation of the dataset. We allocate the first nine years of data (2012 to 2020) to training, keeping 750 randomly drawn surfaces for validation purposes, and use the final year of the dataset (2021) for testing. This yields a training dataset $\mathcal{D}_{\text{train}}$ containing $n_{\text{train}}=43442$ surfaces, a validation dataset $\mathcal{D}_{\text{val}}$ containing $n_{\text{val}}=750$ surfaces and a test dataset $\mathcal{D}_{\text{test}}$ with $n_{\text{test}}=4897$ surfaces.
@@
We remind that we rely on the interpolating graph neural operator introduced in Section 3.2 and described in detail in Appendix B. The model hyperparameters (giving rise to 102529 trainable parameters in total) were identified by manual experimentation and are detailed in Appendix C.2.
@@
We implement a Vega-weighted (see Appendix C.3) version of the fitting loss $\mathcal{L}_{\text{fit}}$ from (8). We compute $\mathcal{L}_{\text{but}}$ directly as (9) (in the transformed coordinates) on a synthetic grid using finite differences. For $\mathcal{L}_{\text{cal}}$, we implement a multiplicative version that is invariant to the level of implied volatility. We provide a precise description in Appendix C.3.
@@
We train the GNO over $500$ epochs on the training dataset using the AdamW optimizer with learning rate $\lambda=10^{-4}$ and weight decay rate $\beta=10^{-5}$, and use a pseudo batch size of 64 by accumulating gradients. We randomly sub-sample the inputs $\mathbf{v}$ and randomize the grids on which we compute the arbitrage losses. The training is performed in around 200 hours using a NVIDIA Quadro RTX 6000 GPU. The validation loss is reported in Appendix C.3 in Table 2. Code and trained model weights are provided in the supplementary material.
@@ 4.2 Results
Figure 3: Benchmark metrics (surface-averages, between 25%- and 75%-quantiles) over training period (computed from $\mathcal{D}_{\text{val}}$) and testing period, as well as surface-averaged spread. Resampled, monthly for training period and weekly for testing period.
Benchmark Metrics
Let $\mathbf{v}=\{v(x)\}_{x\in\pi}$ be the collection of observed implied volatilties and $\hat{v}$ the smoothed surface as produced by a given method. We measure absolute relative error:
$\delta_{\text{abs}}(\hat{v}(x),v(x))=\frac{|\hat{v}(x)-v(x)|}{v(x)}.$
@@
$\delta_{\text{spr}}(\hat{v}(x),v(x))=\frac{2}{s(x)}\big{|}\operatorname{BS}_{\pm}(x,\hat{v}(x))-\operatorname{BS}_{\pm}(x,v(x))\big{|}.$
Since we use simple mid reference prices, $\delta_{\text{spr}}(\hat{v}(x),v(x))\leq 1$ indicates that the prediction $\hat{v}(x)$ for the option $x$ lies within the bid-ask spread.
@@
During production use, the GNO would be retrained regularly using the most recent available data. We emulate this procedure during our evaluation of the benchmark metrics over the test dataset $\mathcal{D}_{\text{test}}$ (containing the year 2021): Following the evaluation of the first month’s test data, the GNO is trained for 10 epochs on the test data just reviewed, with each mini-batch augmented by an equal amount of data from the training dataset $\mathcal{D}_{\text{train}}$. We repeat this, progressively incorporating an additional month of data, until the entire dataset $\mathcal{D}_{\text{test}}$ is assessed. This finetuning-evaluation procedure takes circa 1.8 GPU hours per month.
@@
It is apparent from Figure 3 that our operator deep smoothing approach substantially improves on SVI’s smoothing capabilities, with respect to both $\delta_{\text{abs}}$ and $\delta_{\text{spr}}$. [^7] We produce the SVI benchmark as described in Section E. Our approach, with monthly finetuning, smooths the volatility surface with a MAPE of around 0.5%, while SVI fluctuates between 1% and 2%. The various figures in Appendix D illustrate the qualitative improvements of our method versus SVI. Moreover, our approach appears highly competitive with Ackerer et al., (2020), which performs instance-by-instance volatility smoothing using classical neural networks and reports a MAPE of around 1% for synthetically generated data. [^8] Compare Table 1 of Ackerer et al., (2020). We note that Ackerer et al., (2020) does not perform a similar restriction of the domain of the volatility surface. We reenact its backtesting for the period Jan-Apr 2018 using our method and summarize the results in Table 5 in Appendix C.5.999We emphasize the following aspects. First, the backtest involves dropping half of input datapoints for each surface, and our trained operator continues to perform accurate smoothing, a strong indicator of the robustness of our approach with respect to subsampling of inputs afforded by neural operator’s discretization invariance. Second, while Ackerer et al., (2020) requires to train 61 neural networks to perform the backtest once, our operator approach enables us to run 25 repetitions in around two minutes on a consumer grade laptop CPU, which is the average time that it takes Ackerer et al., (2020) to train one network.
An important fact to consider when analyzing Figure 3 is the historical tightening of the bid-ask spread (displayed in the third row), driven by increases in competition on the S&P 500 option market. This explains why $\delta_{\text{spr}}$ is very small early in the training dataset, both for operator deep smoothing and SVI, while $\delta_{\text{abs}}$ is large: Wide spreads make $\delta_{\text{spr}}$ more lenient an error metric but are accompanied by noisier prices, necessitating greater need for correction by the smoothing algorithm, in turn captured by $\delta_{\text{abs}}$. [^10] This is visually discernible from a comparison of Figure 8 and Figure 10. This argumentation extends to spikes in the spread, which indicate periods of market stress, and then allows to explain spikes in $\delta_{\text{abs}}$.
@@
Figure 4: Average spatial distribution of benchmark metrics and arbitrage terms over $\mathcal{D}_{\text{test}}$.
@@
To test the generalization capabilities of our approach, we procure end-of-day options data for the S&P 500 (SPX), the NASDAQ-100 (NDX), the Dow Jones Industrial Average (DJX), and the Russell 2000 (RUT) for the month of January 2021 from the OptionMetrics Ivy DB US database, accessed by us through the Wharton Research Data Services (WRDS). We evaluate the trained operator on the data for the month of January 2021 (right after the training period on the S&P 500 intraday data), and report the average error metrics $\delta_{\text{abs}}$ and $\delta_{\text{spr}}$ as well as the average arbitrage losses $\mathcal{L}_{\text{cal}}$ and $\mathcal{L}_{\text{but}}$ in Table 1. Firstly, our method maintains its performance on end-of-day S&P 500 data, validating the soundness of our approach: While end-of-day data is slightly different from intraday data, our method still yields small error metrics and arbitrage-free prices. Secondly, the method generalizes well to other indices. We want to stress the fact that our operator has solely been trained on intraday S&P 500 data. Its accurate and virtually arbitrage-free output on end-of-day data of other indices is a strong indicator of the robustness of our approach. We provide further example plots for these datasets in Appendix D.3.
@@
$\delta_{\text{abs}}$ 0.00272 0.01057 0.01629 0.00885
$\delta_{\text{spr}}$ 0.64423 1.53306 0.20736 1.03183

### v3

@@ 3.2 Operator deep smoothing
(a) Smoothed surface with input volatility quotes (blue). $\delta_{\text{abs}}=0.005$, $\delta_{\text{spr}}(\hat{v},\mathbf{v})=1.020$.
(b) Left: Slices of implied volatility scaled by $\sqrt{\tau}$; absence of crossings indicates absence of calendar arbitrage. Right: Implied probability density factor; positivity indicates absence of butterfly arbitrage. In fact, $\mathcal{L}_{\text{cal}}(\theta,\mathbf{v})=0$ and $\mathcal{L}_{\text{but}}(\theta,\mathbf{v})=0$.
Figure 2: Operator deep smoothing of quotes $\mathbf{v}$ from 08.01.2021 at 11:50:00. Compare Figure 13.
Let $\mathbf{v}=\{v(x_{l})\}_{l=1}^{p}$ be the collection of observed data, for example implied volatilities as in Section 2. This notation silently adopts the neural operator philosophy, connecting the data point $v(x_{l})$ with coordinates $x_{l}$ and hinting at a latent function $v\colon D\to\mathbb{R}$ giving rise to the observed values. The smoothing or interpolation task consists of constructing an appropriately regular estimate $\hat{v}$ of $v$ from the data $\mathbf{v}=v|_{\pi}$, with $\pi=\{x_{1},\dots,x_{p}\}$. The operator deep smoothing approach uses a neural operator $\tilde{F}^{\theta}$, trained using historical data, to generate $\hat{v}:=\tilde{F}^{\theta}(\mathbf{v})$. It fundamentally leverages the discretization-invariance to produce consistent results even as the “sensors” $x_{l}$ of the latent function $v$ change in availability and/or location in the domain $D$. This is the situation for volatility smoothing, where the sensors $x_{l}=(\tau_{l},k_{l})$ are the (time-to-expiry, log-moneyness) coordinates of the quoted options, which – as noted in Section 1 and illustrated in Figure 1(b) – move with the market.
@@
Translating the smoothing task to an operator learning problem, we find that the target operator is the (continuous) identity operator
$F_{\text{id}}\colon C(\overline{D})\to C(\overline{D}),\quad v\mapsto v,$
meaning that $F=F_{\text{id}}$ and $\mathcal{A}=\mathcal{U}=C(\overline{D})$. The training dataset is the collection $\mathcal{D}=\{\mathbf{v}^{(i)}\}$ of historical data (labels and features coincide), and we train a suitable neural operator architecture $\tilde{F}^{\theta}$ such that
[…]
where $v$ denotes the postulated latent function of which we observe $\mathbf{v}=v|_{\pi}$, and $\varepsilon$ is some given error tolerance for the smoothing task. Instead of minimizing empirical estimates of the $\infty$-norm $\|\tilde{F}^{\theta}(\mathbf{v})-v\|_{\mathcal{U}}$, however, we suggest a fitting loss based on the root mean square relative error,
@@
for its smoothness and invariance to the scale of the data. [^4] Empirical estimates of $\infty$-norms and $L^{2}$-norms are equivalent loss functions on finite-dimensional spaces. Additional engineering techniques, such as subsampling of inputs during training, are explored in our practical investigation in Section 4.
Constraints
Depending on the application, the smoothing task may be subject to constraints. For volatility smoothing, the smoothed surface $\hat{v}^{\theta}=\tilde{F}^{\theta}(\mathbf{v})$ must be free of static arbitrage (Theorem 2.1). This is effectively enforced by augmenting the loss function with additional penalization terms, moving away from a pure operator learning problem. This does not only promote the relevant properties in the neural operator output but can also help define it when faced with sparsity of data in the domain $D$ (in this context see also Li et al. (2021)). We handle the strike arbitrage constraint 2.1(ii) via
@@
where we ignore the asymptotic condition since our experiments are focused on the bounded domain $D$ (Figure 1(a)). The inclusion of $\varepsilon$ promotes strictly positive implied densities (we will use $\varepsilon=10^{-3}$), while we choose the $1$-norm to induce sparsity in the constraint violation. The calendar arbitrage constraint 2.1(i) can be tackled analogously with
@@
Various neural operator architectures exist, mostly arising from the kernel integral transform framework of Kovachki et al. (2023). Most prominently, these include Fourier neural operators (FNO), delivering state-of-the-art results on fixed grid data, as well as graph neural operators (GNO), able to handle arbitrary mesh geometries (both reviewed in Appendix A.1). While highly effective with documented universality, these neural operators are not directly applicable for interpolation tasks as their layers include a pointwise-applied linear transformation, which limits the output to the set of the input data locations. Dropping this local transformation results in an architecture proved to retain universality Kovachki et al. (2023) and that – at least for its implementation as a GNO – allows to interpolate functions. On the other hand, Lanthaler et al. (2023) proves universality for the architecture combining the local linear transformation with a simple averaging operation, suggesting the fundamental importance of the collaboration of local and non-local components for the expressivity of neural operators. This was noted in Kovachki et al. (2023), for whom retaining the local components can be “beneficial in practice”, and confirmed in our experiments, where a purely non-local architecture led to substantially reduced performance.
@@ 4.1 Model training
We perform our numerical experiments using 20-minute interval CBOE S&P 500 Index Option data from 2012 to 2021. The dataset amounts to a collection of 49089 implied volatility surfaces and just above 60 million individual volatility datapoints (after domain truncation). We refer the reader to Appendix C.1 for details on the preparation of the dataset. We allocate the first nine years of data (2012 to 2020) to training, keeping 750 randomly drawn surfaces for validation purposes, and use the final year of the dataset (2021) for testing. This yields a training dataset $\mathcal{D}_{\text{train}}$ containing $n_{\text{train}}=43442$ surfaces, a validation dataset $\mathcal{D}_{\text{val}}$ containing $n_{\text{val}}=750$ surfaces, and a test dataset $\mathcal{D}_{\text{test}}$ with $n_{\text{test}}=4897$ surfaces.
@@
We employ the interpolating GNO as introduced in Section 3.2 and described in detail in Appendix B. The model hyperparameters (giving rise to 102529 trainable parameters in total) were identified by manual experimentation and are detailed in Appendix C.2. We perform ablations for the connectivity of the graph structure (and thus for the tradeoff between expressivity and computational complexity) in Appendix C.6.
@@
We implement a Vega-weighted version of $\mathcal{L}_{\text{fit}}$ (equation 8). We compute $\mathcal{L}_{\text{but}}$ (equation 9) using finite differences on a synthetic grid. We use a multiplicative formulation of $\mathcal{L}_{\text{cal}}$ that is invariant to the level of implied volatility. We provide a precise description in Appendix C.3 and perform an ablation study for the weighting of the arbitrage losses in Appendix C.7
@@
We train the GNO for $500$ epochs on $\mathcal{D}_{\text{train}}$ using the AdamW optimizer with learning rate $\lambda=10^{-4}$ and weight decay rate $\beta=10^{-5}$, and use a pseudo batch size of 64 by accumulating gradients. We randomly subsample the inputs $\mathbf{v}$ and randomize the grids on which we compute the arbitrage losses. The training is performed in around 250 hours using an NVIDIA Quadro RTX 6000 GPU. The validation loss is reported in Table 2.
@@ 4.2 Results
Figure 3: Benchmark of our method (OpDS) vs. SVI, split between training period (left half; 2012 to 2020) and testing period (right half; 2021). Top and middle: Surface-averages $\langle\delta_{\text{abs}}\rangle$ and $\langle\delta_{\text{spr}}\rangle$; shaded regions depict respective interquartile range. Bottom: Surface-average bid-ask spread $\langle s\rangle$.
Evaluation metrics
Let $\mathbf{v}=\{v(x)\}_{x\in\pi}$ be a collection of observed implied volatilties and $\hat{v}$ the smoothed surface. We measure absolute relative error:
$\delta_{\text{abs}}(\hat{v}(x),v(x))=|\hat{v}(x)-v(x)|/v(x).$
@@
$\delta_{\text{spr}}(\hat{v}(x),v(x))=2|\operatorname{BS}_{\pm}(x,\hat{v}(x))-\operatorname{BS}_{\pm}(x,v(x))|/s(x).$
$\delta_{\text{spr}}(\hat{v}(x),v(x))\leq 1$ indicates that the prediction $\hat{v}(x)$ for the option $x$ lies within the bid-ask spread.
@@
During production use, the GNO would be retrained regularly using the most recent available data. We emulate this procedure for our benchmark in Figure 3. Following the evaluation of the first month of test data (January 2021), the GNO is re-trained for 10 epochs on this data (with each mini-batch augmented by an equal amount of data from the training dataset $\mathcal{D}_{\text{train}}$), before the next month is evaluated. The process is repeated until the entire dataset $\mathcal{D}_{\text{test}}$ is assessed. This finetuning-evaluation procedure takes circa 1.8 GPU hours per month.
Figure 4: Average spatial distribution of benchmark metrics and arbitrage terms over $\mathcal{D}_{\text{test}}$. The terms $\operatorname{But}(\makebox[4.30554pt]{{$\cdot$}},\hat{v},\partial_{k}\hat{v},\partial^{2}_{k}\hat{v})$ and $\partial_{\tau}[\hat{v}\sqrt{\tau}]$ appear in Theorem 2.1. It is apparent that depicted values are strictly positive (indicating absence of arbitrage), lying above the threshold of $\varepsilon=10^{-3}$.
@@
It is apparent from Figure 3 that our operator approach substantially improves on SVI’s smoothing capabilities, with respect to both $\delta_{\text{abs}}$ and $\delta_{\text{spr}}$. [^6] We produce the SVI benchmark as described in Section E. Our approach, with monthly finetuning, smooths the volatility surface with a MAPE of around 0.5%, while SVI fluctuates between 1% and 2%. The various figures in Appendix D illustrate the qualitative improvements of our method over SVI. Moreover, our approach appears highly competitive with Ackerer et al. (2020), which performs instance-by-instance volatility smoothing using classical neural networks and reports a MAPE of around 1% for synthetically generated data. [^7] Compare Table 1 of Ackerer et al. (2020), which we note, however, performs smoothing on a larger strike/expiry domain. We reenact its backtesting for the period Jan-Apr 2018 using our method and summarize the results in Table 5 in Appendix C.5.888We emphasize the following: First, the backtest involves dropping 50% of points for each surface, and our trained operator continues to perform accurate smoothing, a strong indicator of the robustness of our approach with respect to subsampling of inputs. Second, while Ackerer et al. (2020) requires to train 61 neural networks to perform the backtest once, our operator approach enables us to run 25 repetitions in around two minutes on a consumer grade laptop CPU, which is the average time that it takes Ackerer et al. (2020) to train one network.
An important fact to consider when analyzing Figure 3 is the historical tightening of the bid-ask spread $s$ (displayed in the bottom), driven by increased competition on the S&P 500 option market. This explains why $\delta_{\text{spr}}$ is small early in the training dataset, both for our approach and SVI, while $\delta_{\text{abs}}$ is large: Wide spreads make $\delta_{\text{spr}}$ more lenient an error metric but are accompanied by noisier prices, necessitating greater need for correction by the smoothing algorithm, in turn captured by $\delta_{\text{abs}}$. [^9] Compare, e.g., Figure 11 and Figure 13. Similarily, spikes in $s$ (which indicate periods of market stress) help explain spikes in $\delta_{\text{abs}}$.
@@
To test the generalization of our approach, we procure end-of-day options data for the S&P 500 (SPX), the NASDAQ-100 (NDX), the Dow Jones Industrial Average (DJX), and the Russell 2000 (RUT) from the OptionMetrics Ivy DB US database, accessed by us through the Wharton Research Data Services (WRDS). We evaluate the trained operator on the data for the month of January 2021 (right after the training period on the CBOE S&P 500 intraday data), and report the averages of $\langle\delta_{\text{abs}}\rangle$ and $\langle\delta_{\text{spr}}\rangle$ as well as the average arbitrage losses $\mathcal{L}_{\text{cal}}$ and $\mathcal{L}_{\text{but}}$ in Table 1. Firstly, our method maintains its performance on end-of-day S&P 500 data, validating the soundness of our approach: While end-of-day data is slightly different from intraday data, our method still yields small error metrics and arbitrage-free prices. Secondly, the method generalizes well to other indices. We want to stress the fact that our operator has solely been trained on intraday S&P 500 data. Its accurate and virtually arbitrage-free output on end-of-day data of other indices is a strong indicator of the robustness of our approach. We provide further example plots for these datasets in Appendix D.3.
@@
$\langle\delta_{\text{abs}}\rangle$ 0.00272 0.01057 0.01629 0.00885
$\langle\delta_{\text{spr}}\rangle$ 0.64423 1.53306 0.20736 1.03183